OpenStudy (anonymous):

Calculate the average translational kinetic energy of a nitrogen molecule at 27 degree celcius

OpenStudy (anonymous):

Ek=3/2kT k=R/Navocadro Ek= 1.5* 1.4*10^23* (27+273) because T is in kelvin =6.3 10^21 J